OpenStudy (anonymous):

If a number is randomly chosen from the following list, what is the probability that the number is a multiple of 5? [40, 30, 40, 45, 25, 40, 34, 3, 28, 50, 55, 50]

OpenStudy (amistre64):

i see 9 of them that are mults of 5

OpenStudy (a_clan):

9/12

OpenStudy (asadkarim7):

9 /12 3/ 4

OpenStudy (amistre64):

and 12 total gives us: 9/12 = 3/4

OpenStudy (anonymous):

Here's a fun question. In the set of natural numbers from 1 to 19, if 3 numbers are chosen at random and added together, what is the probability that the sum will be a multiple of 3?

OpenStudy (amistre64):

with or without repetition?

OpenStudy (anonymous):

Without.

OpenStudy (amistre64):

20 total in the pool 19+18+17 = 30+15+9 = 40+13 = 53 is the largest sum

OpenStudy (amistre64):

well...54 lol

OpenStudy (amistre64):

which accounts for multiples of 3 up to 18*3

OpenStudy (amistre64):

but since we have 3 chances to pull that pool becomes 20*19*18 in total right?

OpenStudy (amistre64):

6840 ways .... but that doesnt account for duplicates does it lol

OpenStudy (anonymous):

I think you're missing some options.

OpenStudy (amistre64):

im sure I am ;)

OpenStudy (anonymous):

Firstly there's only 19 numbers, not 20.

OpenStudy (anonymous):

1 through 19 inclusive.

OpenStudy (anonymous):

Now you pick 3 of them to sum. The number of possible ways you can pick those 3 is \[{19 \choose 3} = 969\]

OpenStudy (anonymous):

But the real question is, how many will sum to be a multiple of 3?

OpenStudy (anonymous):

To answer this, we have to look at remainders. Each of the numbers can be expressed as: \[n = 3q + r, (q \in \mathbb{N}, r \in \{0,1,2\})\]

OpenStudy (anonymous):

1 for example is 3(0) + 1

OpenStudy (anonymous):

19 is 3(6) + 1

OpenStudy (anonymous):

14 is 3(4)+2

OpenStudy (anonymous):

So we can see that the sum of 3 numbers will be a multiple of 3 if and only if the r's for the three numbers add to a multiple of 3.

OpenStudy (anonymous):

\[n_1 + n_2 + n_3 = 3(q_1 + q_2 + q_3) + r_1 + r_2 + r_3\]

OpenStudy (anonymous):

The possible options for that are: \[r_1 + r_2 + r_3 = 0\]\[r_1 + r_2 + r_3 = 3\]\[r_1 + r_2 + r_3 = 6\] So we can have 0+0+0, 1+2+0, 2+2+2, and 1+1+1

OpenStudy (anonymous):

So we have to go through and enumerate all the remainders in our set. You can make a table: n r 1 1 2 2 3 0 4 1 5 2 6 0 7 1 8 2 9 0 10 1 11 2 12 0 13 1 14 2 15 0 16 1 17 2 18 0 19 1 # of 0's: 6 # of 1's: 7 # of 2's: 6

OpenStudy (amistre64):

lol .... i thought you said it be fun ;)

OpenStudy (anonymous):

it is fun! =)

OpenStudy (anonymous):

So the number of ways we can have (without repetition) 0+0+0 is 6*5*4 = 120 1+1+1 is 7*6*5 = 210 2+2+2 is 6*5*4 = 120 1+2+3 is 6*7*6 = 252 Adding all those possible ways we have 702 options for summing to a multiple of 3. Dividing by the total number of possible ways we can pick 3 numbers we get \[\frac{702}{969} = 72.44\%\]